OpenStudy (anonymous):

how do you find the range and y intercept of y=3sin(2x)+2

jimthompson5910 (jim_thompson5910):

to find the y intercept, plug in x = 0 and evaluate

OpenStudy (anonymous):

y=2?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

the range of sin(x) spans from -1 to +1 in other words, \[\Large -1 \le \sin(x) \le 1\] the x can be replaced with anything you want, so let's replace x with 2x \[\Large -1 \le \sin(2x) \le 1\]

jimthompson5910 (jim_thompson5910):

we can then multiply every side by 3 \[\Large -1*3 \le 3*\sin(2x) \le 3*1\] \[\Large -3 \le 3\sin(2x) \le 3\]

jimthompson5910 (jim_thompson5910):

and then finally add to 2 all sides \[\Large -3+2 \le 3\sin(2x)+2 \le 3+2\] \[\Large -1 \le 3\sin(2x)+2 \le 5\] so the range of y = 3*sin(2x)+2 spans from -1 to +5 (including both endpoints)

OpenStudy (anonymous):

Just to clarify, we must assume always that -1<sin(x)<1?

jimthompson5910 (jim_thompson5910):

yes sin(x) is a function that bounces up and down. It doesn't go past y = 1 or y = -1. It's boxed in so to speak in terms of the y direction

jimthompson5910 (jim_thompson5910):